factorize a^2-6a+9/6a-3
Answer by lenny460(1073)

Factorize:
(a^2 - 6a + 9)/(6a - 3)
a^2 - 6a + 9 = (a - 3)(a - 3) = (a - 3)^2
6a - 3 = 3(2a - 1)
Therefore:
The Answer:
(a - 3)^2 / 3(2a - 1)